1.   You own a stock that will start paying $0.50 annually at the end of the year. It has zero growth in future. If the required rate of return is 14%, what should you pay per share?


Price per share per DDM [Constant dividend formula] = 0.50/14% =3.57

2.   You own a stock that will start paying $0.50 annually at the end of the year. It will then grow each year at a constant annual rate of 5%. If the required rate of return is 14%, what should you pay per share?


Price per share per DDM [Constant growth formula] = 0.50/(14%-5%) = $5.56

3.   What should you pay for a stock assuming you expect the following: a dividend of $1.00 paid at the end of years 1 and 2; cost of equity equal to 8 percent; and, a selling price of $31 at the end of two years?


P0= D1/(1+r) + D2/(1+r)2+ P2/(1+r)2

P0= $1/(1.08) + $1/(1.08)2+ $31/(1.08)2

= $0.9259 + $0.8573 + $26.5775

= $28.36.

 



1
Expert's answer
2021-05-24T15:52:26-0400

1. Price per share using constant dividend formula is:

"P1 = 0.50\/0.14 = 3.57."

2. Price per share is:

"P2 = 0.50\/(0.14 - 0.05) = 5.56."

3. Price per share is:

"P0 = D1\/(1+r) + D2\/(1+r)^2+ P2\/(1+r)^2 = 1\/1.08 + 1\/1.08^2+ 31\/1.08^2\n= 0.9259 + 0.8573 + 26.5775 = 28.36."